OpenStudy (anonymous):

Given segment AB with endpoints A(2x, 4y) and B(10, 22), what is the midpoint?

OpenStudy (whpalmer4):

Midpoint of a line between \((x_1, y_1)\) and \((x_2,y_2)\) is gotten by averaging the corresponding values: \[M = (\frac{1}{2}*(x_1+x_2), \frac{1}{2}*(y_1+y_2))\]

OpenStudy (anonymous):

i got 19?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

No, you should get a point — an x value and a y value.

OpenStudy (whpalmer4):

\[((1/2)(2x + 10), (1/2)(4y + 22))\]

OpenStudy (whpalmer4):

What is \[(1/2)(2x+10)\]? The answer to that will be the x value of the midpoint.

OpenStudy (whpalmer4):

Similarly, the value of \[(1/2)(4y+22)\]will be the y value of the midpoint. Good luck!

OpenStudy (anonymous):

a^2+b^2= c^2 11^2+14^2= c^2, where c is the diagonal. 121+196=c^2 c=sqr(317)= 17.804 or about 18 inches.

OpenStudy (whpalmer4):

No, that's the distance formula...you wanted the midpoint.

OpenStudy (whpalmer4):

Unless you are talking about a new problem now...someone else will have to help, I have to leave.